# Beyond Discretization: Learning the Optimal Solution Path

### v1

@@ 1.1 Our Contributions
The proof of this result utilizes ideas from the convergence of SGD under various “growth conditions” of the gradient (Bottou et al., 2018; Nguyen et al., 2018; Vaswani et al., 2019; Liu et al., 2024; Bertsekas, 1996). See Khaled and Richtárik (2020) for a summary and comparison of these various conditions. Our contribution to this literature is to relate the expressiveness of the basis $\Phi_{1:p}(\lambda)$ to a relaxed, weak-growth condition of Gower et al. (2019, Lemma 2.4). Indeed, we show that under some assumptions, Equation 2 always satisfies this relaxed weak-growth condition, and if the solution path lies in the span of the basis, it satisfies the weak-growth condition of Vaswani et al. (2019). This allows us to leverage results from those works to establish Theorem 3.4.
@@
$\boldsymbol{\Phi}_{1:p}(\cdot)\leftarrow\left[\boldsymbol{\Phi}_{1:{p-1}}(\cdot)\ \boldsymbol{\Phi}_{p}(\cdot)\right]$
@@ 6.1 Weighted Binary Classification
The hyperparameter $\lambda\in[0,1]$ controls the weight placed on each class. We use (scaled and shifted) Legendre polynomials with a $\text{Unif}[0,1]$ distribution.
@@
For LSP, we run SGD using torch.optim.SGD. In lieu of a constant learning rate, we dynamically reduce the learning rate according to the Distance Diagnostic of Pesme et al. (2020, Section 4). This dynamic updating is more reflective of practice. We use the suggested parameters from Pesme et al. (2020) with the exception of $q$, which we tune by examining the performance after $200$ iterations. (We take $q=1.3$) For ALSP, we initialize the algorithm with $5$ polynomials and stop it after adding $12$ polynomials.
Figure 2: LSP: Interpolating the Solution Path for Weighted Binary Classification. Compare $\boldsymbol{\Phi}_{3,1:7}(\lambda)\boldsymbol{\hat{\beta}}$ to $\boldsymbol{\theta}^{\ast}_{3}(\lambda)$ across $\lambda$ after running LSP with $7$ Legendre polynomials for 500, 100, 200, 400 gradient calls.
Figure 1 in the introduction shows the performance for LSP. As we can see, LSP methods converge rapidly to an irreducible error, and this error decreases as we add basis functions. Our adaptive method converges rapidly. This performance is resonated by Figure 2, where we plot the coefficient profile for $\theta^{*}_{3}(\lambda)$ when using $7$ polynomials. The approximation improves as we increase the number of iterations, and eventually become closely aligned with the true solution path.
Practically, this experiment suggests that when the region of $\Lambda$ is compact, taking $\mathbb{P}_{\lambda}$ to be uniform paired with a polynomial basis is favorable. The polynomial basis can then be taken to be orthogonal to $\mathbb{P}_{\lambda}$ as discussed in Section 5.
@@ 6.2 Portfolio Allocation
Unlike our previous experiment, to showcase that the qualitative insights of our theory hold for other algorithms other than SGD, we use torch.optim.BFGS for LSP, ALSP and uniform discretization (for a fair comparison). Unlike SGD, BFGS uses both function and gradient evaluations. We restrict it to use only $10$ function calls per gradient step so that the total work is still proportional to number of gradient calls.
@@ B.1 Calibrating Uniform Discretization
In Figure 4, we compare the solution path error and grid pass error to the target solution path error for our experiments. Figure 4(a) illustrates that in the weighted binary classification experiment with $w(\lambda)=\lambda$, the discretization scheme performs well, satisfying both objectives as the solution path error and grid pass error closely align with the target solution path error. Figure 4(b) demonstrates that for the portfolio allocation problem, both solution path error and grid pass error are still close to each other, but fall below the target solution path error. This is primarily due to the high precision of BFGS, though only very few gradient calls are performed at each grid.
@@ Appendix C Additional Figures
Figure 6: LSP and ALSP for Portfolio Allocation. Compares methods using the first $p=4,9,16$ bivariate-Legendre polynomials. Differs from Figure 3 as the $y$-axis records the number of function evaluations during line-search of BFGS instead of gradient calls.

### v2

@@ 1.1 Our Contributions
The proof of this result utilizes ideas from the convergence of SGD under various “growth conditions” of the gradient (Bottou et al., 2018; Nguyen et al., 2018; Vaswani et al., 2019; Liu et al., 2024; Bertsekas, 1996). See Khaled and Richtárik (2020) for a summary and comparison of these various conditions. Our contribution to this literature is to relate the expressiveness of the basis $\boldsymbol{\Phi}_{1:p}(\lambda)$ to a relaxed, weak-growth condition of Gower et al. (2019, Lemma 2.4). Indeed, we show that under some assumptions, Equation 2 always satisfies this relaxed weak-growth condition, and if the solution path lies in the span of the basis, it satisfies the weak-growth condition of Vaswani et al. (2019). This allows us to leverage results from those works to establish Theorem 3.4.
@@
$\boldsymbol{\Phi}_{1:p}(\cdot)\leftarrow\left[\boldsymbol{\Phi}_{1:{p-1}}(\cdot)\quad\boldsymbol{\Phi}_{p}(\cdot)\right]$
@@ 6.1 Weighted Binary Classification
We consider two different choices of distribution for $\lambda\in[0,1]$ together with two different orthogonal polynomial bases:
  1. i)
    A $\text{Unif}[0,1]$ distribution. Here we use (scaled and shifted) Legendre polynomials.
  2. ii)
[…]
@@
For LSP, we run SGD using torch.optim.SGD. In lieu of a constant learning rate, we dynamically reduce the learning rate according to the Distance Diagnostic of Pesme et al. (2020, Section 4). This dynamic updating is more reflective of practice. We use the suggested parameters from Pesme et al. (2020) with the exception of $q$, which we tune by examining the performance after $200$ iterations. (We take $q=1.3$ for both the Legendre and Jacobi bases.) For ALSP, we initialize the algorithm with $5$ polynomials and stop it after reaching $12$ polynomials ($p=5$ to $12$).
Figure 2: LSP and ALSP for Weighted Binary Classification. (a) Compares methods using Jacobi polynomials. (b) Error in solution path as a function of $\lambda$ using Jacobi polynomials. (c) and (d) compare $\boldsymbol{\theta}_{3}(\lambda)$ across $\lambda$ in the Legendre and Jacobi bases respectively with $7$ polynomials.
Figure 1 in the introduction and Figure 2 panel (a) show the performance for the Legendre and Jacobi bases respectively. As we can see, LSP methods converge rapidly to an irreducible error, and this error decreases as we add basis functions. Our adaptive method converges rapidly. This performance is resonated by Figure 2 panels (c) and (d), where we plot the coefficient profile for $\theta^{*}_{3}(\lambda)$ when using $7$ Legendre and Jacobi polynomials. The approximation improves as we increase the number of iterations, and eventually become closely aligned with the true solution path.
Figure 2 panel (b) shows that the largest errors occur almost periodically at a few places across $\Lambda=[0,1]$, matching the prediction by the equioscillation theorem.
Practically, this experiment suggests that when the region of $\Lambda$ is compact, taking a polynomial basis paired with a distribution supported everywhere over $\Lambda$ is favorable. The polynomial basis can then be taken to be orthogonal to $\mathbb{P}_{\lambda}$ as discussed in Section 5.
@@ 6.2 Portfolio Allocation
Unlike our previous experiment, to showcase that the qualitative insights of our theory hold for other algorithms other than SGD, we use torch.optim.LBFGS for LSP, ALSP and uniform discretization (for a fair comparison). Unlike SGD, L-BFGS uses both function and gradient evaluations. We restrict it to use only $10$ function calls per gradient step so that the total work is still proportional to the number of gradient calls.
@@ B.1 Calibrating Uniform Discretization
In Figure 4, we compare the solution path error and grid pass error to the target solution path error for our experiments. Figure 4(a) illustrates that in the weighted binary classification experiment, the discretization scheme performs well, satisfying both objectives as the solution path error and grid pass error closely align with the target solution path error. Figure 4(b) demonstrates that for the portfolio allocation problem, both solution path error and grid pass error are still close to each other, but fall below the target solution path error. This is primarily due to the high precision of L-BFGS, though only very few gradient calls are performed at each grid.
@@ Appendix C Additional Figures
Figure 6: LSP and ALSP for Portfolio Allocation. Compares methods using the first $p=4,9,16$ bivariate-Legendre polynomials. Differs from Figure 3 as the $y$-axis records the number of function evaluations during line-search of L-BFGS instead of gradient calls.